OpenStudy (anonymous):

another brain freeze.. ln4x=14 ?

OpenStudy (anonymous):

how do i solve this?

OpenStudy (sriram):

e^14=4x x=e^14 ------ 4

OpenStudy (anonymous):

\[\ln(4x)=14\] \[e^{14}=4x\] \[x=\frac{e^{14}}{4}\]

OpenStudy (anonymous):

X=300651.1(corr to 1d.p)